OpenStudy (anonymous):

What is the sum of the first 19 terms of an arithmetic series with a rate of increase of 7 and a7 = 46? A)1,197 B)1,273 C)1,373 D)1,423 E)1,327

OpenStudy (anonymous):

sum of n numbers Sn = n/2[2a + (n – 1)d] a7= a + 6d a + 6 × 7 = 46 a = 46 – 42 = 4 a is first term = 4 d = common difference = 7 n = total numbers = 19 S10 = 19/2[8 + 18 × 7 = 19/2[114 = 1273 --------- B

OpenStudy (anonymous):

THANKS FOR THE MEDAL

OpenStudy (anonymous):

no thank you <3

OpenStudy (anonymous):

Jk

OpenStudy (anonymous):

lol